# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v6

@@ V-C Implementation Details
To benchmark against existing methods[8, 31, 17, 18, 19, 34, 35, 35, 11, 39, 32, 21], we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
V-D Quantitative Results
V-D1 Results on FoggyLane
@@
We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods. As shown in the Tabel II, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
@@
From the Table II, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
V-D2 Results on FoggyCULane
@@
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table III, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
V-D3 Results on FoggyTusimple
@@
As shown in the Table IV, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods. Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%. This further demonstrates the effectiveness of our method in foggy conditions.
TABLE IV: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN[8] AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD[11] ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD[11] ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT[17] CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT[17] CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT[17] CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR[34] WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA[31] AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA[31] AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane[39] ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane[39] ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane[39] ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet[18] CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet[18] CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet[18] CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet[19] IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet[19] IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet[35] IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet[35] IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane[32] IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN[21] IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN[21] IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
V-E Ablation Study
@@
We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone. As shown in Tables V and VI, each module improves the model’s performance. The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information. Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane. Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
TABLE V: Ablation experiments of different components on the FoggyLane dataset.
@@
TABLE VI: Ablation experiments of different components on the FoggyCULane dataset.

### v7

@@ 5.3 Implementation Details
To benchmark against existing methods(Pan et al., 2018; Zheng et al., 2021; Tabelini et al., 2021a; Zheng et al., 2022; Ran et al., 2023; Liu et al., 2021b; Dai et al., 2024; Qin et al., 2020; Liu et al., 2021a; Wang et al., 2025), we followed the training settings provided in the respective papers. For FoggyCULane and FoggyTusimple, we adopted the training parameters from the original CULane and Tusimple datasets, while for FoggyLane, we extended the training epochs to ensure reliable results.
5.4 Quantitative Results
5.4.1 Results on FoggyLane
@@
We evaluated our proposed method on the FoggyLane dataset and compared it with other leading lane detection methods. As shown in the Tabel 2, we mainly compared the methods in terms of F1 scores, computational complexity (GFlops), processing speed (FPS).
@@
From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
5.4.2 Results on FoggyCULane
@@
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
5.4.3 Results on FoggyTusimple
@@
As shown in the Table 4, the traffic scenes in the FoggyTusimple dataset are relatively simple, resulting in smaller performance differences between methods. Nonetheless, our method achieved an F1 score of 96.95, outperforming the best existing method by 0.43%. This further demonstrates the effectiveness of our method in foggy conditions.
5.5 Ablation Study
@@
We conducted ablation studies on the FoggyLane and FoggyCULane datasets to evaluate the effectiveness of the GFFM, KFFM, and LEEM modules, using Swin-t as the backbone. As shown in Tables 5 and 6, each module improves the model’s performance. The GFFM module contributes the most, increasing the F1-score by 1.07 on FoggyLane and 0.41 on FoggyCULane, highlighting its ability to capture both local and global information. Adding the KFFM module further boosts the F1-score by 0.63 on FoggyLane and 0.26 on FoggyCULane. Finally, the LEEM module improves edge detection, increasing the F1-score by 0.28 on FoggyLane and 0.14 on FoggyCULane.
Table 5: Ablation experiments of different components on the FoggyLane dataset.
@@
Table 6: Ablation experiments of different components on the FoggyCULane dataset.